# CLAP: Isolating Content from Style through Contrastive Learning with Augmented Prompts

### v5

@@ 5.3 More Analysis
Prompt Augmentation Combinations
We explore diverse combinations of our tailored prompt augmentation techniques and examine the impact of Easy Data Augmentation (EDA) methods [42] on the VLCS dataset. All evaluated techniques demonstrate that CLAP provides consistent improvements over CLIP. Further details are presented in § A.2.
Prompt Sources
We assess the influence of different training prompt sources on performance, using the VLCS dataset and incorporating EDA techniques. The evaluated prompt formats include: (1) our template-based prompts; (2) LLM-generated prompts from ChatGPT-3.5 [3], with generation details in § B.2; (3) prompts of the form "a [random] style of [class]," where "[random]" is sampled from a random word generator [^1] https://github.com/vaibhavsingh97/random-word; and (4) prompts generated by the PromptStyler method [9]. The results suggest that training with simpler, more structured prompts tends to yield better performance. Detailed quantitative results are reported in § A.3.
Experiments on Different Model Scales
We repeat our zero-shot evaluations using larger pre-trained CLIP models, specifically ViT-L/14 and ResNet50x16. Across both model scales, CLAP consistently improves zero-shot performance while also reducing variance. These findings reaffirm CLAP’s scalability and robustness in enhancing CLIP representations. Quantitative results are provided in §§ B.3 and B.4.
@@ 6 Conclusion
To enhance the representations of pre-trained CLIP-like models, this study investigates the disentanglement of latent content variables. Adopting a causal perspective on the generative processes underlying vision-language data, we demonstrate that training a disentangled network in one modality enables effective disentanglement across both modalities. Leveraging the semantic richness and logical structure of text data, we find that disentanglement is more readily achieved in the language modality through targeted text augmentation.
Based on these insights, we propose CLAP (Contrastive Learning with Augmented Prompts), a method designed to acquire disentangled vision-language content features. Extensive experiments validate the effectiveness of CLAP, showing substantial improvements in zero-shot and few-shot classification tasks, along with enhanced robustness to input perturbations. We hope this work will motivate further research into disentangled representation learning within multimodal frameworks.
Appendix A More on Experiments with ViT-B/16
A.1 Implementation Details
This section provides detailed implementation settings for our experiments using the CLIP pre-trained ViT-B/16 model.
@@ A.1 Implementation Details
The network’s output dimension is set to match the 512-dimensional CLIP feature space, thereby eliminating the need for feature downsampling. The latent dimension is tailored for each dataset: 256 for PACS, 448 for OfficeHome, and 512 for VLCS and DomainNet, to accommodate differences in category counts and dataset complexity. The weight parameter $\alpha$ is adjusted per dataset as follows: 0.208 for PACS, 0.056 for VLCS, 0.14 for OfficeHome, and 0.2 for DomainNet. During training, $\alpha$ is consistently set to 1.
@@
We use consistent training parameters across all datasets. The Adam optimizer is employed with a learning rate of 0.0001. Training is limited to 8,000 steps, with the average loss evaluated every 480 steps, and early stopping applied if the loss does not decrease by at least 0.01 over five checkpoints. Batch sizes are set to 8 for PACS and VLCS, 96 for OfficeHome, and 384 for DomainNet. The temperature parameter $\tau$ is configured to 0.5 for PACS and VLCS, and 0.3 for OfficeHome and DomainNet. The loss coefficient $\lambda$ is set to 1 for PACS and VLCS, and 0.0001 for OfficeHome and DomainNet, reflecting the fewer number of classes in the former two datasets.
Prompt augmentations OSD, OCD, and SPO are applied across all datasets with a probability of 0.5. For PACS and VLCS, Gaussian noise (mean 0, standard deviation 0.02) is randomly inserted at the beginning, middle, or end of the augmented prompts to diversify the training samples. For linear probe evaluations in the few-shot setting, we use L2 normalization and cross-entropy loss, training over 1,000 epochs with a batch size of 32. Early stopping is used with a patience of 10 epochs and a loss improvement threshold of 0.001.
@@
We train a disentangled network using image augmentations and the InfoNCE loss with a temperature parameter $\tau$ set to 0.5. Image augmentations include random cropping (scale $\in[0.64,1.0]$) and color distortion (brightness = 0.5, hue = 0.3), each applied with a probability of 0.5. All other training and inference configurations for Im.Aug are consistent with those used for CLAP across the datasets.
A.2 Prompt Augmentation Combinations
In table 6, we evaluate various combinations of our tailored prompt augmentation techniques alongside Easy Data Augmentation (EDA) [42] on the VLCS dataset. Across all configurations, CLAP consistently enhances CLIP’s performance while effectively reducing performance variance across different prompt conditions.
Among the tested combinations, OSD+OCD+SPO+IGN (①②⑤⑥) yields the highest average zero-shot accuracy and the lowest variance, outperforming all other augmentation strategies, including EDA. Notably, even without the inclusion of random noise (IGN), CLAP achieves substantial improvements over CLIP, particularly in mitigating performance degradation under noisy prompt conditions, as reflected by a significantly lower $\Delta_{(NC)}$.
Table 6: Effect of prompt augmentation combinations on VLCS. We assess the impact of various augmentations: OSD (①), OCD (②), ITD (③), ASD (④), SPO (⑤), and IGN (⑥). ZS(Avg.) denotes the average zero-shot accuracy across four different inference prompts. CLAP improves both accuracy and robustness, with the best results achieved using the combination ①②⑤⑥.
Metric CLIP (base) EDA ①②③④⑤⑥ ①②③④⑤ ①②③④⑥ ①②③④ ③④⑤ ①②⑤⑥
@@ A.2 Prompt Augmentation Combinations
A.3 Prompt Sources
In table 7, we evaluate the impact of different training prompt formats, sourced from various synthetic origins, on the performance of CLAP using the VLCS dataset. All experiments in this subsection are conducted with EDA (Easy Data Augmentation) techniques applied during training.
The evaluated prompt formats include: Template, representing our structured template-based prompts; LLM, referring to prompts generated using ChatGPT-3.5 [3] (see § B.2 for generation details); Random, which uses the format "a [random] style of [class]", where "[random]" is filled using a random word generator; and Prm.Stl., which denotes vectorized prompts produced by the PromptStyler method [9].
Table 7: Effect of training prompt sources on CLAP performance (VLCS). We train CLAP using EDA-augmented prompts from various sources. All prompt types improve CLIP’s baseline zero-shot performance. Simpler prompt formats—Random and Template—achieve the highest accuracy and robustness.
Metric CLIP (base) LLM Random Prm.Stl. Template
@@ A.3 Prompt Sources
The results indicate that CLAP trained with any of the evaluated prompt sources leads to consistent improvements over the original CLIP model. Interestingly, despite the sophistication of LLM-generated and PromptStyler-based prompts, the simpler Random and Template prompts outperform them in both accuracy and robustness.
It is worth noting, however, that while these alternative prompt formats paired with EDA enhance zero-shot performance, they do not surpass the performance achieved by prompt augmentations specifically optimized for the template-based format (as reported in § A.2).
A.4 Detailed Results on ViT-B/16
Details on Zero-Shot Evaluations
We present domain-level zero-shot performance using various prompts across each dataset in table 8. CLAP consistently improves upon CLIP across all prompt variants and domains. Notably, as CLAP exclusively uses text data for training, it retains CLIP’s domain generalization capabilities—an outcome of CLIP’s large-scale training corpus. Rather than compromising this generalization, CLAP enhances it through more effective disentanglement of content information, leading to consistent zero-shot gains across all evaluated domains.
Table 8: Domain-level zero-shot results with ViT-B/16. We report top-1 accuracy (%) using four different prompts: ZS(C), ZS(CP), ZS(PC), and ZS(NC), across dataset domains. CLAP improves zero-shot performance in all settings.
Datasets Domain-level avg. top-1 acc. (%) of zero-shot performance usig ViT-B/16 (↑)
@@ A.4 Detailed Results on ViT-B/16
Table 9: Domain-level few-shot results with ViT-B/16. We report top-1 accuracy (%) for 1-shot through 32-shot settings across dataset domains. CLAP outperforms both CLIP and Im.Aug, particularly in low-shot scenarios.
Datasets Domain-level avg. top-1 acc. (%) of few-shot performance of ViT-B/16 (↑)
@@
Details on Few-Shot Evaluations
We present detailed quantitative results of few-shot performance in table 9. CLAP consistently enhances few-shot capabilities, demonstrating improvements across all evaluated domains and test datasets.
Details on Adversarial Evaluations
In table 10, we report domain-level adversarial performance for PACS, VLCS, OfficeHome, and DomainNet. CLAP improves both zero-shot and one-shot robustness across all domains. Although Im.Aug strengthens one-shot robustness, its impact on zero-shot adversarial performance remains inconsistent.
Details on Ablative Analysis
table 11 presents the detailed results of zero-shot performance using various combinations of prompt augmentation strategies. Additionally, table 12 summarizes the effects of key hyperparameters—including temperature $\tau$, latent dimension size, and weight coefficient $\alpha$—evaluated across individual domains. These findings confirm that CLAP maintains effectiveness across a broad range of hyperparameter configurations.
Table 10: Domain-level adversarial results (ViT-B/16). Average top-1 accuracy (%) under adversarial attacks: FGSM, PGD-20, and CW-20, evaluated in both zero-shot (ZS-C) and one-shot settings across domains of PACS, VLCS, OfficeHome, and DomainNet.
Datasets FGSM PGD-20 CW-20
@@
PACS A 76.3 79.3 79.3 61.2 78.0 87.3 1.7 2.2 1.8 16.0 42.1 63.1 1.5 2.0 2.3 0.5 1.1 1.7
C 94.9 95.0 94.0 66.5 84.2 95.1 33.3 37.7 35.6 33.3 57.2 86.1 28.8 34.0 33.2 11.9 23.6 31.8
P 91.6 90.3 91.7 67.4 80.8 92.1 5.7 7.0 6.7 27.1 55.0 69.8 4.7 4.9 5.8 0.7 2.7 4.1
S 84.5 87.5 89.8 71.6 74.6 83.8 75.8 78.4 79.2 63.0 66.3 74.6 74.5 76.8 77.9 62.7 65.4 70.3
VLCS C 55.3 53.8 55.5 25.8 28.8 25.3 4.4 5.1 4.7 2.0 5.2 2.5 2.9 3.1 3.5 0.7 1.2 1.0
L 49.4 45.5 50.6 27.0 32.6 30.4 15.2 14.9 16.0 6.4 8.9 8.0 12.4 11.2 13.0 6.1 8.3 7.7
S 61.7 58.1 62.5 48.0 46.9 51.6 13.2 13.9 14.0 8.6 10.7 10.0 9.2 8.8 10.2 8.3 7.9 8.4
V 65.3 63.2 65.6 36.5 40.1 41.0 7.5 7.9 7.9 5.3 9.4 8.9 5.2 4.8 5.6 2.9 2.8 2.9
O.H. A 55.3 53.8 55.5 25.8 28.8 25.3 4.4 5.1 4.7 2.0 5.2 2.5 2.9 3.1 3.5 0.7 1.2 1.0
C 49.4 45.5 50.6 27.0 32.6 30.4 15.2 14.9 16.0 6.4 8.9 8.0 12.4 11.2 13.0 6.1 8.3 7.7
P 61.7 58.1 62.5 48.0 46.9 51.6 13.2 13.9 14.0 8.6 10.7 10.0 9.2 8.8 10.2 8.3 7.9 8.4
R 65.3 63.2 65.6 36.5 40.1 41.0 7.5 7.9 7.9 5.3 9.4 8.9 5.2 4.8 5.6 2.9 2.8 2.9
D.N. C 57.8 50.9 58.8 33.3 34.3 35.0 21.6 18.7 22.8 18.4 19.6 20.0 15.8 12.5 16.6 7.0 7.5 7.8
I 35.8 28.0 37.0 12.2 13.3 13.2 6.1 3.7 6.7 4.6 5.3 5.1 3.3 1.9 3.7 0.9 0.9 0.9
P 43.9 39.0 44.3 18.4 20.6 20.3 3.1 2.8 3.3 8.6 10.4 9.9 1.8 1.3 1.9 0.3 0.3 0.3
Q 12.9 10.3 13.2 10.9 10.8 11.1 8.4 6.8 8.6 5.4 5.4 5.6 7.1 5.4 7.4 4.9 4.8 5.1
R 62.1 55.9 62.4 34.5 35.9 36.5 7.1 6.5 7.5 17.6 19.7 19.6 4.5 3.4 4.7 1.2 1.4 1.4
S 49.1 43.3 49.7 25.7 26.0 27.5 17.8 15.5 18.6 13.6 14.4 15.1 13.4 10.2 13.9 5.0 5.2 5.6
Appendix B Data Synthesis
B.1 Synthetic Image Generation

### v6

@@ 5.3.2 Ablations
5.3.3 Prompt Augmentation Combinations
We explore diverse combinations of our tailored prompt augmentation methods and examine Easy Data Augmentation (EDA) techniques [42] on the VLCS dataset. Each tested technique showcases CLAP’s enhancements over CLIP, with details available in Sec. 0.A.2.
5.3.4 Prompt Sources
We assess the impact of different training prompt formats, originating from various synthetic sources, on the performance of the VLCS dataset, incorporating EDA techniques. Our evaluation includes our template-based prompts, LLM-generated prompts by ChatGPT-3.5 [3] (with the generation process detailed in Sec. 0.B.2 ), prompts structured as "a [random] style of [class]," where "[random]" is filled with terms from a random word generator [^2] https://github.com/vaibhavsingh97/random-word, and prompts produced using the PromptStyler method [9]. The findings indicate that the training prompts with simpler forms tend to yield better performance, with detailed quantitative results presented in Sec. 0.A.3.
5.3.5 Experiments on Different Model Scales
In our repeated experiments assessing zero-shot performance on the ViT-L/14 and ResNet50x16 pre-trained with CLIP, we consistently find that CLAP improves zero-shot performance while also reducing performance variances. This consistent observation underscores CLAP’s effectiveness in enhancing the quality of CLIP representations. For quantitative details supporting these findings, please see the Appendix 0.C.
@@ 6 Conclusion
To enhance pre-trained CLIP-like models, this study delves into disentangling latent content variables. Through a causal analysis of the underlying generative processes of vision-language data, we discover that training a disentangled network in one modality can effectively disentangle content across both modalities. Given the high semantic nature of text data, we identify that disentanglement is more achievable within the language modality through text augmentation interventions. Building on these insights, we introduce CLAP (Contrastive Learning with Augmented Prompts) to acquire disentangled vision-language content features. Comprehensive experiments validate CLAP’s effectiveness, demonstrating significant improvements in zero-shot and few-shot performance, and enhancing robustness against perturbations. We anticipate that our work will inspire further exploration into disentangling latent variables within vision-language models.
Appendix 0.A More on Experiments with ViT-B/16
0.A.1 Implementation Details
In this section, we detail the implementation of our experiments utilizing the CLIP pre-trained ViT-B/16 model:
@@ 0.A.1 Implementation Details
The network’s output dimension is aligned with the 512-dimensional CLIP features, thereby obviating the need for input feature downsampling. The latent dimensions are tailored to each dataset: 256 for PACS, 448 for OfficeHome, and 512 for VLCS and DomainNet, to accommodate the variety of categories and complexity of datasets. The weight parameter $\alpha$ is adjusted to 0.208 for PACS, 0.056 for VLCS, 0.14 for OfficeHome, and 0.2 for DomainNet, while it is consistently maintained at 1 throughout the training phase.
@@
Training parameters are consistent across datasets, employing the Adam optimizer with a learning rate of 0.0001, limiting training to 8,000 steps with checking the average loss every 480 steps, and instituting early stopping after five checkpoints without a loss decrease of at least 0.01. Batch sizes are adjusted to 8 for PACS and VLCS, 96 for OfficeHome, and 384 for DomainNet, with the temperature parameter $\tau$ set at 0.5 for PACS and VLCS, and 0.3 for OfficeHome and DomainNet. The loss coefficient $\lambda$ is set to 1 for PACS and VLCS, and 0.0001 for OfficeHome and DomainNet, due to the first two datasets have less classes. Prompt augmentations, OSD+OCD+SPO, are applied across datasets all with a 0.5 probability. For the PACS and VLCS datasets, Gaussian noise with a zero mean and a standard deviation of 0.02 is randomly inserted at the beginning, middle, or end of the augmented-view prompts to enrich the training samples. In the linear probe evaluations for few-shot analysis, L2 normalization and cross-entropy loss are utilized for training over 1,000 epochs with a batch size of 32, incorporating early stopping with a patience threshold of 10 epochs and a loss decrease criterion of 0.001.
@@
We train a disentangled network using image augmentation, applying the InfoNCE loss with a temperature parameter $\tau$ set to 0.5. This include image augmentation techniques, image cropping ($scale\in[0.64,1.0]$) and color distortion ($brightness=0.5,hue=0.3$), each with a probability of 0.5. Other training and inference configurations for Im.Aug are consistent with those used for CLAP across all datasets.
0.A.2 Prompt Augmentation Combinations
In Tab. 6, we explore different combinations of our tailored prompt augmentation techniques and EDA (Easy Data Augmentation) [42] techniques on the VLCS dataset. Each combination demonstrates CLAP’s effectiveness in enhancing CLIP’s performance and reducing performance disparities. The combination of OSD+OCS+SPO+IGN achieves the highest average accuracy and the least variance, outperforming the EDA techniques. Notably, even without incorporating random noise in the augmentations, CLAP significantly surpasses CLIP in handling perturbations on prompts, as evidenced by the largely reduced $\Delta_{(NC)}$.
Table 6: We evaluate prompt augmentation combinations on the VLCS dataset: OSD (①), OCD (②), ITD (③), ASD (④), SPO (⑤), and IGN (⑥). ZS(Avg.) shows average zero-shot accuracy acoss four distinct inference prompts. CLAP boosts CLIP’s accuracy and reduces variances, with ①②⑤⑥ as the optimal combination.
Metrics CLIP (base) Avg. top-1 acc. (%) of different augmentations
EDA ①②③ ①②③ ①②③ ①②③ ③④⑤ ①②⑤
④⑤⑥ ④⑤ ④⑥
@@ 0.A.2 Prompt Augmentation Combinations
0.A.3 Prompt Sources
In Tab. 7, we examine the effects of various training prompt formats, sourced from different synthetic origins, on the VLCS dataset performance, utilizing
Table 7: We employ EDA augmentation to train CLAP with diverse prompt sources on the VLCS dataset. Each prompt source contributes to improvements in CLIP’s zero-shot performance, with "Random" and "Template" prompts, in their simpler forms, yielding better outcomes.
Metrics CLIP (base) Avg. top-1 acc. (%) of different sources
  LLM  Random Prm.Stl. Template
@@ 0.A.3 Prompt Sources
EDA techniques. The prompt formats are defined as follows: "Template" refers to the template-based prompts fundamental to our primary approach; "LLM" designates prompts created by ChatGPT-3.5 [3], with the generation process elaborated in Sec. 0.B.2; "Random" describes prompts formatted as "a [random] style of [class]," with "[random]" being replaced by terms from a random word generator; and "Prm.Stl." indicates vectorized prompts generated through PromptStyler[9].
Our experimental results indicate that CLAP, when trained across these varied prompt formats, enhances the performance of CLIP. Notably, despite the complex generation mechanisms of "LLM" and "Prm.Stl." prompts, the simpler, random-styled and template-based prompts demonstrate superior efficacy. However, it is important to highlight that the improvements attributed to these diverse prompt formats, trained with EDA, do not surpass the best performance of the prompt augmentations tailored for template-based prompts.
0.A.4 Detailed Results on ViT-B/16
0.A.4.1 Details on Zero-Shot Evaluations
We present the domain-level zero-shot performance with various prompts across each dataset in Tab. 8. CLAP consistently enhances CLIP’s zero-shot performance across these different prompts. Given that CLAP exclusively utilizes text data for training, it does not compromise CLIP’s inherent ability to generalize across domains, which is acquired from its extensive training dataset. Rather, by achieving a more effective disentanglement of content, it unequivocally enhances CLIP’s zero-shot performance across all dataset domains.
Table 8: Domain-level zero-shot results of the ViT-B/16 model on the test datasets.
Dataset Domains Domain-level avg. top-1 acc. (%) of zero-shot performance usig ViT-B/16 (↑)
@@ 0.A.4.1 Details on Zero-Shot Evaluations
Table 9: Domain-level few-shot results of the ViT-B/16 model using the test datasets.
Dataset Domains Domain-level avg. top-1 acc. (%) of few-shot performance of ViT-B/16 (↑)
@@
0.A.4.2 Details on Few-Shot Evaluations
We display the quantitative results of few-shot performance in Tab. 9. CLAP consistently enhances the few-shot capabilities, showcasing improvements across test datasets at a closer domain level.
0.A.4.3 Details on Adversarial Evaluations
In Tab. 10, we detail our adversarial performance evaluations for PACS, VLCS, OfficeHome, and DomainNet, respectively. CLAP enhances both zero-shot and one-shot performance across all domains of the tested datasets. While Im.Aug boosts one-shot robustness against adversarial tasks, its impact on zero-shot adversarial robustness is inconsistent.
0.A.4.4 Details on Ablative Analysis
In Tab. 11, we provide detailed results from our analysis on zero-shot performance using various combinations of prompt augmentations. Additionally, in Tab. 12, we present the outcomes of our ablative studies focusing on the hyperparameters $\tau$, latent dimension, and $\alpha$, respectively, each evaluated domain-wise. The results indicate that CLAP is effective across a wide range of hyperparameters.
Table 10: Domain-level results under adversarial attacks of ViT-B/16 on the datasets.
Dataset Domains Domain-level avg. top-1 acc. (%) under adversarial attackings using ViT-B/16 (↑)
FGSM PGD-20 CW-20
@@ 0.A.4.4 Details on Ablative Analysis
PACS A. 76.3 79.3 79.3 61.2 78.0 87.3 1.7 2.2 1.8 16.0 42.1 63.1 1.5 2.0 2.3 0.5 1.1 1.7
C. 94.9 95.0 94.0 66.5 84.2 95.1 33.3 37.7 35.6 33.3 57.2 86.1 28.8 34.0 33.2 11.9 23.6 31.8
P. 91.6 90.3 91.7 67.4 80.8 92.1 5.7 7.0 6.7 27.1 55.0 69.8 4.7 4.9 5.8 0.7 2.7 4.1
S. 84.5 87.5 89.8 71.6 74.6 83.8 75.8 78.4 79.2 63.0 66.3 74.6 74.5 76.8 77.9 62.7 65.4 70.3
VLCS C. 55.3 53.8 55.5 25.8 28.8 25.3 4.4 5.1 4.7 2.0 5.2 2.5 2.9 3.1 3.5 0.7 1.2 1.0
L. 49.4 45.5 50.6 27.0 32.6 30.4 15.2 14.9 16.0 6.4 8.9 8.0 12.4 11.2 13.0 6.1 8.3 7.7
S. 61.7 58.1 62.5 48.0 46.9 51.6 13.2 13.9 14.0 8.6 10.7 10.0 9.2 8.8 10.2 8.3 7.9 8.4
V. 65.3 63.2 65.6 36.5 40.1 41.0 7.5 7.9 7.9 5.3 9.4 8.9 5.2 4.8 5.6 2.9 2.8 2.9
OfficeHome A. 55.3 53.8 55.5 25.8 28.8 25.3 4.4 5.1 4.7 2.0 5.2 2.5 2.9 3.1 3.5 0.7 1.2 1.0
C. 49.4 45.5 50.6 27.0 32.6 30.4 15.2 14.9 16.0 6.4 8.9 8.0 12.4 11.2 13.0 6.1 8.3 7.7
P. 61.7 58.1 62.5 48.0 46.9 51.6 13.2 13.9 14.0 8.6 10.7 10.0 9.2 8.8 10.2 8.3 7.9 8.4
R. 65.3 63.2 65.6 36.5 40.1 41.0 7.5 7.9 7.9 5.3 9.4 8.9 5.2 4.8 5.6 2.9 2.8 2.9
DomainNet C. 57.8 50.9 58.8 33.3 34.3 35.0 21.6 18.7 22.8 18.4 19.6 20.0 15.8 12.5 16.6 7.0 7.5 7.8
I. 35.8 28.0 37.0 12.2 13.3 13.2 6.1 3.7 6.7 4.6 5.3 5.1 3.3 1.9 3.7 0.9 0.9 0.9
P. 43.9 39.0 44.3 18.4 20.6 20.3 3.1 2.8 3.3 8.6 10.4 9.9 1.8 1.3 1.9 0.3 0.3 0.3
Q. 12.9 10.3 13.2 10.9 10.8 11.1 8.4 6.8 8.6 5.4 5.4 5.6 7.1 5.4 7.4 4.9 4.8 5.1
R. 62.1 55.9 62.4 34.5 35.9 36.5 7.1 6.5 7.5 17.6 19.7 19.6 4.5 3.4 4.7 1.2 1.4 1.4
S. 49.1 43.3 49.7 25.7 26.0 27.5 17.8 15.5 18.6 13.6 14.4 15.1 13.4 10.2 13.9 5.0 5.2 5.6
Appendix 0.B Data Synthesis
0.B.1 Synthetic Image Generation